Determine the type of bond in each pair of atoms. Which is the most polar bond? Least polar? Support your answers by calculating the electronegativity difference in each pair.






Cl - F, O - S, P - N, Na - Cl

1
Expert's answer
2022-04-08T10:47:04-0400

Cl – F – covalent polar, ΔEN = 3.98 – 3.16 = 0.82 > 0.4   

O – S – covalent polar, ΔEN = 3.44 – 2.58 = 0.86 > 0.4

P – N – covalent polar, ΔEN = 3.04 – 2.19 = 0.85 > 0.4  

Na – Cl – ionic bond between cation Na+ and anion Cl-, ΔEN = 3.16 – 0.93 = 2.23 >> 0.4

The most polar is Na – Cl, the least polar is Cl – F.